OpenStudy (anonymous):

i need help on how to solve factors, I'm in geometry! please help me, i have a D in this class!

OpenStudy (anonymous):

which factor?

OpenStudy (anonymous):

its this 4x+20x+120=0

OpenStudy (anonymous):

\[4x^2+20x+120=0\] or what?

OpenStudy (anonymous):

yeah what you wrote, how do you solve it?

OpenStudy (across):

It's helpful to first simplify the expression into a form that looks something like this:\[x^2+ax+b=0,\]where \(a\) and \(b\) are numbers (if you can).

OpenStudy (anonymous):

First, you factor out the 4 because all 3 coefficients are divisible by 4. \[4(x^2+5x+30)=0\] Which you can simplify to \[x^2+5x+30=0\] Then you use the quadratic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] where a=1, b=4, and c=30

OpenStudy (vishal_kothari):

roots are imaginary.....

OpenStudy (vishal_kothari):

i think b=5....

OpenStudy (anonymous):

wait what? whats the quadratic formula? my teacher hasn't thought us that yet.

OpenStudy (anonymous):

@sammyiko: Time to sue your teacher.

OpenStudy (across):

The quadratic formula is a means to come up with the roots of a polynomial of the form\[ax^2+bx+c=0,\]\[\forall a, b, c \in \mathbb{R}.\]

OpenStudy (anonymous):

Also, if b^2-4ac is negative, then there is no solution to the problem

OpenStudy (anonymous):

If you are in geometry then your teacher probably wants you solve it graphically ...

OpenStudy (anonymous):

so in this case, there is no solution because 5^2-4*1*30=25-120 which is negative

OpenStudy (turingtest):

it's not really NO solution, just imaginary

OpenStudy (anonymous):

Also, if b^2-4ac is negative, then there is no solution to the problem --- You mean no real solutions isn't ?

OpenStudy (anonymous):

just imaginary

OpenStudy (anonymous):

i feel so stupid! i have no idea what you said, okay so its x2(squared)+5x+30=0 so its (x+5)(x+6)?

OpenStudy (anonymous):

WHAAAAT!

OpenStudy (turingtest):

if you haven't studied imaginary numbers yet that's ok

OpenStudy (across):

No because\[(x+5)(x+6)=x^2+11x+30.\]

OpenStudy (anonymous):

When you graph it, it doesn't cross the x-axis, so there isn't a solution because the solutions to a quadratic equation are where it crosses the x-axis.

OpenStudy (anonymous):

so no solution to 4x2(squared)+20x+25=0?

OpenStudy (turingtest):

It's not a zero on the real plane, but if you allow complex ones...

OpenStudy (anonymous):

I've found this website useful when I'm lazy http://www.wolframalpha.com/

OpenStudy (anonymous):

and no, there isn't a solution if you're in high school geometry.

OpenStudy (anonymous):

4x2+20x+120=0

OpenStudy (anonymous):

yeah I'm in high school geometry

OpenStudy (turingtest):

true, not for high school geometry. I think the numbers should stay real, especially if you don't know imaginary ones.

OpenStudy (anonymous):

i went to that website, but i need to know how to solve this, my teacher said its like algebra 2 trig stuff.

OpenStudy (turingtest):

trig? are you supposed to know trig?

OpenStudy (anonymous):

Really? I learned about quadratic equations in Algebra I

OpenStudy (anonymous):

no! but he said we're suppose to learn a little, he's crazy

OpenStudy (anonymous):

You can just show him the graph of the function

OpenStudy (anonymous):

is this right? x2-16x+55 is (x+5)(x-11)?

OpenStudy (turingtest):

no... (x+5)(x-11)=x^2-6x-55

OpenStudy (anonymous):

sorry...x^2-16x+55 isn't factorable, I don't think

OpenStudy (anonymous):

wait whats ^ suppose to mean?

OpenStudy (turingtest):

to the power of x^2 is x(squared)

OpenStudy (anonymous):

oh okay so its not factorable

OpenStudy (turingtest):

can you factor x^2-x-2 ???

OpenStudy (turingtest):

BTW x^2-16x+55=(x-5)(x-11)

OpenStudy (anonymous):

oh okay! what about 6x^2+29x+35? i don't get how to solve the 6 thingy

OpenStudy (turingtest):

these are ugly... no common factors...

OpenStudy (anonymous):

so its "not factorable"?

OpenStudy (turingtest):

It is, it's just a tricky one are you very comfortable with things like x^2-x-2 ??? factor if you can

OpenStudy (anonymous):

um ye(:

OpenStudy (turingtest):

are you sure? it's pointless to do more difficult ones if you can't do this what's the answer?

OpenStudy (anonymous):

(x-2)(x+1)?

OpenStudy (turingtest):

good! so now I guess we'll try 6x^2+29x+35 There is a more efficient way to do this, but here's how I do itWe know there will be + signs all through because the last number is positive. what numbers should be in the blanks? on the right of each set of parentheses they will nee do be numbers that multiply to 35, so what can they possibly be?